# RL Zero: Zero-Shot Language to Behaviors without any supervision

### v1

@@ C.1 Zero-shot imitation: Discriminator vs Discriminator-free
C.2 Cross Embodiment experiments
Table 4 describes the videos used for cross-embodiment along with the win rate of the behaviors generated by RLZero when compared to a base model which trains SMODICE (Ma et al., 2022a) on the nearest states found with the same grounding methods as RLZero.
Prompt Descriptions Video Link/Meta AI Prompt Win rate vs SMODICE
human in backflip position animated human trying backflip 2/5
@@ C.2 Cross Embodiment experiments
cow yoga pose Moves (2019) 5/5
downward facing dog yoga pose with one leg raised in the air Moves (2019) 5/5
lying on back with one leg raised in the air Nicole (2021) 5/5
lying on back with both legs raised in the air LivestrongWoman (2014) 5/5
high plank yoga pose Well+Good (2019) 5/5
sitting down with legs laid in the front Calisthenicmovement (2021) 5/5
bending forward while standing on one leg with the other leg raised in the back Yoga (2022) 3/5
front splits where both legs stretched out along the same line Suarez (2022) 4/5
Table 4: Comparison of Win rates vs SMODICE
C.3 Imagination-Free RLZero Complete Results
@@
C.4 More Failure Cases

### v2

@@ C.2 Zero-shot imitation: Discriminator vs Discriminator-free
C.3 Cross Embodiment experiments
Tables  C.3 and  C.3 describe the videos used for cross-embodiment along with the win rate of the behaviors generated by RLZero when compared to a base model which trains SMODICE [46] on the nearest states found with the same grounding methods as RLZero.
{NiceTabular}
c|c|c|c Prompt Descriptions Video Link/Meta AI Prompt Win rate vs SMODICE
Stickman (2D Humanoid) Human in backflip position animated human trying backflip 2/5
@@ C.3 Cross Embodiment experiments
Cow yoga pose [51] 5/5
Downward dog with one leg raised in the air [51] 5/5
Lying on back with one leg raised in the air [53] 5/5
Lying on back with both legs raised in the air [41] 5/5
High plank yoga pose [81] 5/5
Sitting down with legs laid in the front [11] 5/5
Warrior III pose [87] 3/5
Front splits [70] 4/5
Table 5: Comparison of Win rates vs SMODICE for Stickman
{NiceTabular}
c|c|c|c Prompt Descriptions Video Link/Meta AI Prompt Win rate vs SMODICE
SMPL (3D Humanoid)
A karate kick position a karate kick 5/5
A cat doing a handstand a side profile of cat doing headstand 5/5
An arabesque ballet position ballet movement 5/5
Animated wikiHow demo of a cartwheel [83] 5/5
Running running 5/5
Lying crunches [82] 5/5
Plank position [81] 5/5
Table 6: Comparison of Win rates vs SMODICE for 3D SMPL Humanoid
C.4 Imagination-Free RLZero Complete Results
@@
C.5 More Failure Cases
@@
C.6 RLZero evaluation with Video-Embedding similarity
In this section, we experiment with another metric for comparison – embedding similarity between a video of the generated behavior and the text. We use InternVideo2 to embed the videos and take the cosine similarity with the prompt used to generate the behavior. Table C.6 shows the results for this metric of comparison. Unfortunately, we observed that the similarity score is frequently higher even for behaviors that differ significantly from the prompt. This points to a limitation of using this metric for evaluation. Some reasons for this failure could be the limited context length of 8 for the video embedding model or a misalignment between video and text embedding vectors [40].
{NiceTabular}
c|cc|cc|c Image-language reward   Video-language reward   RLZero
IQL TD3 (Base Model) TD3 IQL
Walker
Quadruped
Cheetah
Stickman
[…]